can someone help? doing question from my lab report and im confused on finding the answer to this. molar mass used for everything else was 39.997g/mol
4. If you assume that each pellet of sodium hydroxide is 100mg, how many moles of sodium hydroxide are in 3 pellets?
Answer #1

1) Mass of 1 pellet = 100 mg = 100 / 1000 = 0.1 g

Therefore, mass of 3 pellets is 3 x 0.1 g = 0.3 g.

2) Moles of NaOH in 3 pellets = Mass(g) / MW

= 0.3 g / 39.997 g/mol

= 0.00750 mole.

= 7.50 x 10-3 mole.

Molesof NaOH = 7.50 x 10-3 moles.